You can do one half of a job in an hour your friend can do one third of the same job in an hour how long will it take to do the job if you work together
Answer by fractalier(6550)

That means you can do the whole job in two hours and he can do it in three. The set up then becomes
t/2 + t/3 = 1
Multiply by 6 and get
3t + 2t = 6
5t = 6
t = 6/5 = 1.2 hours = 1 hour 12 minutes